OpenStudy (anonymous):

solve -3x^2+6x-2=3x-8 by factoring I guess

OpenStudy (anonymous):

solve -3x^2+6x-2=3x-8 add 2 to both sides -3x^2+6x-2+2 =3x-8 + 2 -3x^2+6x =3x-6 -3x(x-2) = 3(x-2) cancel out (x-2) -3x=3 divide both sides by -3 x=-1

OpenStudy (anonymous):

thanks.